How much heat, in joules is required to heat one nail from 29 C to its melting point? Express your answer to three significant figures and include the appropriate units.


1
Expert's answer
2021-02-11T04:46:16-0500

Q = m×c×"\Delta T"

Q = heat required

m = mass of one nail = we take average mass of one nail = 1g

c = specific heat of iron = 0.452 J/gC

melting point of iron =1535⁰C

"\Delta T" = 1535 - 29 = 1506⁰C

Q = 1×0.452×1506

Q = 680.71 J